# An Information-Theoretic Framework for Receiver Quantization in Communication

### v1

@@ Appendix D Proof of Lemma 7
=\;= $\mathcal{A}-\sum\limits_{k=0}^{K-1}\ell\cdot\left(k^{2}\ell^{2}\cdot\exp\frac{-k^{2}\ell^{2}}{2}\right)$ (167b)
=\;= $\mathcal{A}-\mathcal{C}$ (167c)
@@
=\;= $\sqrt{\frac{2}{\pi}}\mathcal{B}-\sum\limits_{k=0}^{K-1}\ell\cdot\left(k^{2}\ell^{2}\exp\frac{-k^{2}\ell^{2}}{2}\right)$ (168b)
=\;= $\sqrt{\frac{2}{\pi}}\mathcal{B}-\mathcal{C},$ (168c)
@@ Appendix E Proof of Proposition 11
First, (80) and (81) can be obtained by the Pythagorean relations (73) and (74), respectively. We begin from (24c) and (25) and rewrite them as
@@
$\mathrm{E}\left[|\mathsf{Y}|^{2}\right]=\mathsf{E}\left[|\mathsf{V}|^{2}\right]\frac{2}{\pi}\mathcal{B},$ (170)
@@
$\mathrm{E}\left[\mathsf{V}\overline{\mathsf{Y}}\right]=(\mathcal{E}_{x}+\sigma^{2})(1-\mathsf{mmse})=\mathrm{E}\left[|\mathsf{Y}|^{2}\right].$ (172)
Substituting (172) into (171) yields
@@
thereby implying that (52) is equivalent to (77). Combine (171) and its equivalent form (173), we obtain

### v2

@@ Appendix D Proof of Lemma 7
$=\;$ $\mathcal{A}-\sum\limits_{k=0}^{K-1}\ell\cdot\left(k^{2}\ell^{2}\cdot\exp\frac{-k^{2}\ell^{2}}{2}\right)$ (167b)
$=\;$ $\mathcal{A}-\mathcal{C}$ (167c)
@@
$=\;$ $\sqrt{\frac{2}{\pi}}\mathcal{B}-\sum\limits_{k=0}^{K-1}\ell\cdot\left(k^{2}\ell^{2}\exp\frac{-k^{2}\ell^{2}}{2}\right)$ (168b)
$=\;$ $\sqrt{\frac{2}{\pi}}\mathcal{B}-\mathcal{C},$ (168c)
@@ Appendix E Proof of Proposition 11
First, (79) and (80) can be obtained by the Pythagorean relations (72) and (73), respectively. According to the definitions of $\mathsf{X}$, $\mathsf{Y}$, and $\mathsf{V}$, we can rewrite (26c) and (27) as
@@
$\mathrm{E}\left[|\mathsf{Y}|^{2}\right]=\frac{|h|^{2}\sigma_{x}^{2}+\sigma^{2}}{2}\mathrm{E}\left[|Y|^{2}\right]=\mathsf{E}\left[|\mathsf{V}|^{2}\right]\frac{2}{\pi}\mathcal{B},$ (170)
@@
$\mathrm{E}\left[\mathsf{V}\overline{\mathsf{Y}}\right]=(\mathcal{E}_{x}+\sigma^{2})(1-\mathsf{mmse})=\mathrm{E}\left[|\mathsf{Y}|^{2}\right],$ (172)
where the second equality follows from the length of $\mathbf{Y}$; see the Pythagorean relation (73). Substituting (172) into (171) and utilizing the fact $\mathrm{E}[|\mathsf{X}|^{2}]=\mathrm{E}[\mathsf{X}\overline{\mathsf{V}}]$ yield
@@
thereby implying that (51) is equivalent to (77). Combining (171), its equivalent form (173), and the definition of the Pearson correlation coefficient (18), we obtain